Two systems K and K synchronize their clocks at t = t = 0 when their origins are aligned as system k passes by system k al
Answer #1

Let k in stationary and kin moving at v= 0.8200 rclative to k As observed from the K reference fram : 7 The proton gun is fiI-V2 (2 - 10.2 c)2 0.82CX 1 C2 C2 -0.6924 Vo. 3276 tarta - var) I where zu - x - VZ) x2 = 2 (M2 - V22) t = 1.747 (3n (3nc- =